x + y = 3k
x - y = k
The solution of the system shown is _____.

Answer by Edwin McCravy(20086)

x + y = 3k
x - y = k


Add the equations term by term just as they are and the terms in y will cancel:

 x + y = 3k
 x - y =  k
-----------
2x     = 4k

Divide both sides by 2

     x = 2k


Multiply the second equation x - y = k through by -1 making it
-x + y = -k

Then add it to the first equation term by term and the terms in x will cancel:

 x + y = 3k
-x + y = -k
-----------
    2y = 2k

Divide both sides by 2

     y = k

Answer (x,y) = (2k,k)

Edwin